Question 4) Consider an economy with two kinds of firms :

  • Firms with sticky prices have ??

  • Firms with flexible prices ??
    A) Derive the following SRAS that we discussed in the class. You have to explain your steps and

    assumption clearly ( 10 points)

?=?̅ + α (P – EP),
Where α = s/[(1-s)a] , and
s=0.5 is the fraction of firms with sticky prices;
a=0.5 is just a parameter.
?̅= 200
B) How do you interpret the SRAS? ( 5 points)
C) Assume that the AD curve is given as Y= 1000 – 10P, and we are in the long-run equilibrium. Determine the values of Y, P and EP in the equilibrium. (15 points)

D) Suppose that the share of firms with sticky prices increase to s=0.75. What will be the impact of that on your answer in part b? Calculate it? (10 points)
Answer #1

A. A) Firms set their desired prices as: Pf = P + a(y-7) ps = Ep + a (EY-EF) for sticky price firms, they expect to output to0.75 D) Let S : 0.75, :. aa (i-S).a ^ 0.25 x 0.5 - 6 :: SRAS: Y = 200 + 6 (P-EP) However as LRÁS: Y: 200 and P : EP with AD: